OpenStudy (anonymous):

if a printer takes 8 hours to print class schedules for all the students in college. a faster printer can do the job in 6 hours. how long will it take to do the job if both printers are used?

OpenStudy (shubhamsrg):

momma ?

OpenStudy (anonymous):

You use rates: So: 1/8 + 1/6 = 1/x, x = 24/7 hours for both printers: So about 3 hours and 25 minutes.

OpenStudy (anonymous):

thanks